The perimeter around a rectangle is 48 inches. The length of the rectangle is 2 inches more than the width. What is the width of the rectangle?
Answer by rfer(16322)

2x+2(x+2)=48
2x+2x+4=48
4x=44
x=11 in
x+2=13 in